how would you do this? Determine whether quadrilateral ABCD with the given vertices is a parallologram. Explain.
A(2,5), B(5,9), C(6,3), D(3,-1)


problem 2:
A(-1,6), B(2, -3), C(5,9), D(2,7)

Answer by fcabanski(1391)

Remember that the slopes of parallel lines are equal. Find the slope of the sides and make sure the opposite sides have the same slopes.


You will have to graph it to see that AB and CD are opposite sides, and AD and BC are opposite sides. The coordinates are (x,y) and the slope of each side is (y2-y1)/(x2-x1).


AB and CD must have the same slopes. For AB it's (9-5)/(5-2) = 4/3 and for CD it's (-1-3)/(3-6) = -4/-3 = 4/3. Those sides are parallel.


AD and BC must have the same slopes. For AD it's (5-(-1))/(2-3) = 6/-1 = -6 and for BC it's (9-3)/(5-6) = -6. That figure is a parallelogram because the opposite sides are parallel.


Graph 2. It's not a parallelogram.

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com